Cindy is working her way through school. She works two part-time jobs for a total of 24 hours a week. Job A pays $5.70 per hour, and Job B pays $6.60 per hour. How many hours did she work at each job the week that she made $150.30.
Found 2 solutions by poliphob3.14, mananth:
Answer by poliphob3.14(115)

Solution:Let x hours that Cindy makes $6.6 per hour, at job B
Then 24-x hours that Cindy makes $5.7 per hour, at job A
Since she made $150.3 per week, we write the equation.
6.6x+5.7(24-x)=150.3 => 6.6x-5.7x=150.3-5.7*24 = > 0.9x=8.1 => x=9.
Answer;Cindy work 9 hours at job B and 24-9=15 hours at job A.

Answer by mananth(16949)

Job I number of hours ----------------x
Job II number of hours ----------- y
x+y=24 .............1
5.7x+6.6y=150.3 .............2
Eliminate y
multiply (1)by -6.6
Multiply (2) by 1
-6.6x-6.6y=-158.4
5.7x+6.6y=150.3

Add the two equations
-0.9x=-8.1

/-0.9

x=9 Job I number of hours

plug value of x in (1)

x+y=24
9+y=24
y=24-9
y=15 Job II number of hours